# Physics-Constrained Fine-Tuning of Flow-Matching Models for Generation and Inverse Problems

### v1

@@ Abstract
We present a framework for fine-tuning flow-matching generative models to enforce physical constraints and solve inverse problems in scientific systems. Starting from a model trained on low-fidelity or observational data, we apply a differentiable post-training procedure that minimizes weak-form residuals of governing partial differential equations (PDEs), promoting physical consistency and adherence to boundary conditions without distorting the underlying learned distribution. To infer unknown physical inputs, such as source terms, material parameters, or boundary data, we augment the generative process with a learnable latent parameter predictor and propose a joint optimization strategy. The resulting model produces physically valid field solutions alongside plausible estimates of hidden parameters, effectively addressing ill-posed inverse problems in a data-driven yet physics-aware manner. We validate our method on canonical PDE benchmarks, demonstrating improved satisfaction of PDE constraints and accurate recovery of latent coefficients. Our approach bridges generative modelling and scientific inference, opening new avenues for simulation-augmented discovery and data-efficient modelling of physical systems.
@@ 1 Introduction
Physical systems with rich spatio-temporal structure can be effectively represented by deep generative models, including diffusion and flow-matching methods (Kerrigan, Migliorini, and Smyth 2023; Erichson et al. 2025; Baldan et al. 2025; Price et al. 2024). Although their dynamics can be highly complex, these systems are often governed by fundamental principles, such as conservation laws, symmetries, and boundary conditions, that constrain the space of admissible solutions. Incorporating such physical structure into generative modelling can improve both sample fidelity and out-of-distribution generalization.
@@
In many scientific domains, including atmospheric and oceanographic modelling, seismic inversion, and medical imaging, we often observe system states without access to the underlying physical parameters that govern them. Crucially, PDE-based constraints are typically parameter-dependent, with residuals that vary according to material properties, source terms, or other latent variables. Prior work has largely focused on simple or global constraints—such as fixed boundaries or symmetries, that apply uniformly across the data distribution. Handling parameter-dependent constraints naively would require training over the joint distribution of solutions and parameters, which is often infeasible because parametric labels are missing, expensive to obtain, or high-dimensional.
Addressing this limitation is critical for scientific discovery. Many inverse problems in the natural sciences and engineering require reasoning about unobserved parameters or exploring hypothetical scenarios inaccessible to direct experimentation. A generative model that can enforce parameter-dependent PDE constraints using only observational data would provide a powerful tool for data-efficient simulation, hypothesis testing, and the discovery of new physical phenomena, helping to bridge the gap between raw observations and mechanistic understanding.
@@
This work proposes a framework for fine-tuning flow-matching generative models to enforce parameter-dependent PDE constraints without requiring joint parameter–solution training data. This work aligns with a growing trend of simulation-augmented machine learning (Karniadakis et al. 2021), where generative models accelerate scientific discovery by efficiently exploring physically plausible solution spaces. Our approach reformulates fine-tuning as a stochastic optimal control problem via Adjoint Matching (Domingo-Enrich et al. 2025), guided by weak-form PDE residuals. By augmenting the model with a latent parameter evolution, we enable joint generation of physically consistent solution–parameter pairs, addressing ill-posed inverse problems.
We evaluate our proposed fine-tuning framework on two canonical physical systems. We demonstrate denoising and conditional generation capabilities, including robustness to noisy data and the ability to infer latent parameters from sparse observations. Visual and quantitative results, including strong reductions in residuals across tasks and robustness to model misspecification, highlight the flexibility of our method for integrating physical constraints into generative modelling.
@@ 2 Related Work
Integrating physical constraints—such as boundary conditions, symmetry invariances, and partial differential equation (PDE) constraints—into machine learning models improves both accuracy and out-of-distribution generalization. Classical approaches, such as Physics-Informed Neural Networks (PINNs) (Raissi, Perdikaris, and Karniadakis 2019), directly regress solutions that satisfy governing equations. While effective for forward or inverse problems, PINNs do not capture distributions over solutions, making them unsuitable for generative tasks that require sampling diverse plausible outcomes.
In the generative setting, the main challenge is ensuring that the physically constrained samples retain the variability of the underlying generative model, avoiding pathalogical issues such as mode collapse. Bastek, Sun, and Kochmann (2024) proposes a unified framework for introducing physical constraints into Denoising Diffusion Probabilistic Models (DDPMs) (Ho, Jain, and Abbeel 2020) at pre-training time, by adding a first-principles physics-residual loss to the diffusion training objective. This loss penalises violations of governing PDEs (e.g. fluid dynamics equations) so that generated samples inherently satisfy physical laws. The method was empirically shown to reduce residual errors for individual samples significantly, while simultaneously acting as a regulariser against overfitting, thereby improving generalisation. To evaluate the physics-residual loss, one needs to compute the expected PDE residual of the final denoised sample conditioned on the current noisy state in the DDPM process. Accurately estimating this expectation requires generating multiple reverse-diffusion trajectories from the same noisy sample, which makes pre-training significantly more expensive. A common alternative is to use Tweedie’s formula to approximate the conditional expectation in a single pass, but this shortcut introduces bias, particularly in the final denoising steps.
Zhang and Zou (2025) proposes enforcing constraints through a post-hoc distillation stage, where a deterministic student model is trained from a vanilla diffusion model to generate samples in one-step, regularised by a PDE residual loss. In (Wang et al. 2025) the authors introduce PhyDA, diffusion-based data assimilation framework that ensures reconstructions obey PDE-based dynamics, specifically for atmospheric science. An autoencoder is used to encode sparse observations into a structured latent prior for the diffusion model, which is trained with an additional physical residual loss.
@@
Various works have proposed approaches to enforce PDE constraints at inference time, often in combination with observational constraints, drawing connections to conditional diffusion models (Dhariwal and Nichol 2021; Ho and Salimans 2021). Huang et al. (2024) introduce guidance terms within the denoising update of a score-based diffusion model to steer the denoising process towards solutions which are both consistent with data and underlying PDEs. A related approach was considered by Xu et al. (2025), further introducing an adaptive constraint to mitigate instabilities in early diffusion steps. In (Christopher, Baek, and Fioretto 2024), the authors recast the inference-time sampling of a diffusion process as a constrained optimization problem, each diffusion step is projected to satisfy user-defined constraints or physical principles. This allows strict enforcement of hard constraints (including convex and non-convex constraints, as well as ODE-based physical laws) on the generated data. Lu and Xu (2024) consider the setting where the base diffusion model is trained on cheap, low-fidelity simulations, leveraging a similar approach to generate down-scaled samples via projection.
@@
Flow-matching (FM) (Lipman et al. 2023) has emerged as a flexible generative modelling paradigm for complex physical systems across science, including molecular systems (Hassan et al. 2024), weather (Price et al. 2024) and geology (Zhang, Li, and Huang 2024) . In the context of physics-constrained generative models (Utkarsh et al. 2025) introduces a zero-shot inference framework to enforce hard physical constraints in pretrained flow models, by repeatedly projecting the generative flow at sampling time. Similarly, (Cheng et al. 2024) proposed the ECI algorithm, to adapt a pretrained flow-matching model so that it exactly satisfies constraints without using analytical gradients. In each iteration of flow sampling, ECI performs: an Extrapolation step (advancing along the learned flow), a Correction step (applying a constraint-enforcement operation), and an Interpolation step (adjusting back towards the model’s trajectory). While projection approaches are a compelling strategy for hard constraints, they can be challenging particularly for local constraints such as boundary conditions, as direct enforcement can introduce discontinuities. The above approach mitigates this by interleaving projections with flow steps, however this relies on the flow’s ability to rapidly correct such non-physical artifacts.
Baldan et al. (2025) propose Physics-Based Flow Matching (PBFM), which embeds constraints (PDE or symmetries) directly into the FM loss during training. The approach leverages temporal unrolling to refine noise-free final state predictions and jointly minimizes generative and physics-based losses without manual hyperparameter tuning of their tradeoff. To mitigate conflicts between physical constraints and the data loss, they employ the ConFIG (Liu, Chu, and Thuerey 2024), which combines the gradients of both losses in a way that ensures that gradient updates always minimise both losses simultaneously.
@@
Related to our approach are the works on generative models for Bayesian inverse problems (Stuart 2010), where the goal is to infer distributions over latent PDE parameters given partial or noisy observations. Conditional diffusion and flow-matching models can be used to generate samples from conditional distributions and posterior distributions, supporting amortized inference and uncertainty quantification (Song et al. 2021; Utkarsh et al. 2025; Zhang, Rao, and Agrawala 2023). Conditioning is typically achieved either through explicit parameter inputs or guidance mechanisms during sampling, as in classifier-guided diffusion. While effective when large volumes of paired training data is available, these approaches are less relevant to observational settings where parameters are unobserved. In contrast, our approach connects the observed data to the latent parameters only during post-training, requiring substantially smaller volumes of data.
@@ 3 Method
FM models are trained to learn and sample from a given distribution of data $p_{\text{data}}$. They approximate this distribution by constructing a Markovian transformation from noise to data, such that the time marginals of this transformation match those of a reference flow $X_{t}=\beta_{t}X_{1}+\gamma_{t}X_{0}$. Specifically FM models learn a vector field $v_{t}(x)$ that transports noise to data, via the ODE $dX_{t}=v_{t}(X_{t})$. We can optionally inject a noise schedule $\sigma(t)$ along the trajectory to define an equivalent SDE that preserves the same time marginals (Maoutsa, Reich, and Opper 2020),
$\begin{split}dX_{t}&=\left(v_{t}(X_{t})+\frac{\sigma(t)^{2}}{2\eta_{t}}\left(v_{t}(X_{t})-\frac{\dot{\beta}_{t}}{\beta_{t}}X_{t}\right)\right)+\sigma(t)\,dB_{t}\\[3.0pt] &=:b_{t}(X_{t})+\sigma(t)\,dB_{t},\end{split}$
where $\eta_{t}=\gamma_{t}\bigl{(}\tfrac{\dot{\beta}_{t}}{{\beta}}\gamma_{t}-\dot{\gamma}_{t}\bigr{)}$.
@@
To achieve this, we leverage the adjoint-matching framework of (Domingo-Enrich et al. 2025). This work reformulates reward fine-tuning for flow-based generative models as a control problem in which the base generative process given by $v_{t}^{\text{base}}$ is steered toward high-reward samples via modifying the learned vector field, which we denote as $v_{t}^{\text{ft}}$ with corresponding drift term $b_{t}^{\text{ft}}$. Our approach is conceptually related to reward- or preference-based fine-tuning of generative models (Christiano et al. 2017; Sun et al. 2024), where a learned or computed reward steers generation toward desired properties. Here, the reward is defined via PDE residuals, encoding knowledge about underlying dynamics and physical constraints to the solutions space as deviations to differential operators or boundary conditions.
@@
Reward
A generative model can reproduce the visual characteristics of empirical data while ignoring the physics that governs it, thereby rendering the samples unusable for downstream scientific tasks. To bridge this gap we impose the known governing equations as soft constraints, expressed through differential operators $\mathcal{L}_{\alpha}x=0$ with parameters $\alpha$. Throughout, a generated sample $x$ is interpreted as the discretisation of a continuous field $x(\xi)$ on a domain $\Omega$. Boundary information is captured either by Dirichlet conditions $x(\xi)=h(\xi)$ or Neumann conditions $\partial_{n}x(\xi)=\tilde{h}(\xi)$ for $\xi\in\partial\Omega$.
The strong PDE residual is defined as
$\mathcal{R}_{\text{strong}}(x,\alpha)=\bigl{\|}\mathcal{L}_{\alpha}x\bigr{\|}_{L^{2}(\Omega)}^{2},$
and deviations from the boundary data are measured by $\mathcal{R}_{\text{BC}}$. In practice, strong residuals involve high-order derivatives that make the optimization landscape unstable. We therefore adopt weak-form residuals, which only require first-order derivatives and remain numerically stable under noisy or misspecified data, of the form $\langle\mathcal{L}_{\alpha}x,\psi\rangle_{L^{2}(\Omega)}$ for suitably chosen test functions $\psi\in\Psi$. Repeated applications of integration-by-parts will transfer derivatives from $x$ to $\psi$ until only first‑order derivatives of $x$ remain, which is sufficient for the problems considered.
The set $\Psi$ is composed of radially symmetric, compactly supported local polynomial kernels. For each evaluation we draw $N_{\text{test}}$ such functions; their centers and length‑scales are sampled at random. A mollifier enforces $\psi|_{\partial\Omega}=0$, justifying the integration by parts. The resulting residual is
@@
These randomly sampled local test functions act as stochastic probes of PDE violations, providing a low-variance, data-efficient learning signal. A more detailed description of the test functions used can be found in Appendix C.
Joint Evolution
@@
$\hat{x}_{1}\;=\;x_{t}+(1-t)\,v_{t}^{\text{base}}(x_{t}),\qquad\hat{\alpha}_{1}\;=\;\varphi(\hat{x}_{1})$
The direction from the current state $\alpha_{t}$ to the predicted final parameter $\hat{\alpha}_{1}$ serves as a base vector field which we use to evolve alpha, i.e. $v_{t,\,\alpha}^{\text{base}}(\alpha_{t})=(\hat{\alpha}_{1}-\alpha_{t})/(1-t)$ inducing corresponding drift $b_{t,\alpha}^{\text{base}}$. This surrogate base flow, starting at a noise sample $\alpha_{0}^{\text{base}}\sim\mathcal{N}(0,I)$, emulates a denoising process of the recovered parameter. We denote by $\alpha^{\text{base}}$ the parameter aligned with the base trajectory $x^{\text{base}}$. While the evolution of $\alpha^{\text{base}}$ does not influence the trajectory of $x^{\text{base}}$, the inferred vector field can be used to effectively regularize the generation of the fine-tuned model.
Similarly, to regularise towards the parameter recovered under the base model, we introduce an additional field $v_{t,\,\alpha}^{\text{reg}}(\alpha_{t}^{\text{ft}})=(\hat{\alpha}_{1}^{\text{base}}-\alpha_{t}^{\text{ft}})/(1-t)$. This vector field points from the current parameter estimate of the fine-tuned trajectory $\alpha_{t}^{\text{ft}}$ to the recovered parameter under the base model $\hat{\alpha}_{1}^{\text{base}}$. The field is used to pull the fine‑tuned dynamics towards final samples associated with parameters similar to those of the base trajectory. The introduced vector fields are visualized in Fig. 1.
Figure 2: Denoising results for the Darcy test case for the same noise samples. Left: noisy input pressure field and its inferred permeability. Middle: fine-tuned output with parameter regularisation $(\lambda_{f}=0.1)$. Right: fine-tuning without regularisation. With regularisation the fine-tuned solution effectively suppresses noise while preserving the permeability-induced structural details captured by the base model. (Color maps throughout this work taken from (Crameri, Shephard, and Heron 2020))
Adjoint Matching

### v2

@@ Abstract
We present a framework for fine-tuning flow-matching generative models to enforce physical constraints and solve inverse problems in scientific systems. Starting from a model trained on low-fidelity or observational data, we apply a differentiable post-training procedure that minimizes weak-form residuals of governing partial differential equations (PDEs), promoting physical consistency and adherence to boundary conditions without distorting the underlying learned distribution. To infer unknown physical inputs, such as source terms, material parameters, or boundary data, we augment the generative process with a learnable latent parameter predictor and propose a joint optimization strategy. The resulting model produces physically valid field solutions alongside plausible estimates of hidden parameters, effectively addressing ill-posed inverse problems in a data-driven yet physics-aware manner. We validate our method on canonical PDE problems, demonstrating improved satisfaction of physical constraints and accurate recovery of latent coefficients. Further, we confirm cross-domain utility through fine-tuning of natural-image models. Our approach bridges generative modelling and scientific inference, opening new avenues for simulation-augmented discovery and data-efficient modelling of physical systems.
@@ 1 Introduction
Physical systems with rich spatio-temporal structure can be effectively represented by deep generative models, including diffusion and flow-matching methods (Kerrigan et al., 2024; Erichson et al., 2025; Baldan et al., 2025; Price et al., 2023). Although their dynamics can be highly complex, these systems are often governed by fundamental principles, such as conservation laws, symmetries, and boundary conditions, that constrain the space of admissible solutions. Incorporating such physical structure into generative modelling can improve both sample fidelity and out-of-distribution generalization.
@@
In many scientific domains, including atmospheric and oceanographic modelling, seismic inversion, and medical imaging, we often observe system states without access to the underlying physical parameters that govern them. Crucially, PDE-based constraints are typically parameter-dependent, with residuals that vary according to material properties, source terms, or other latent variables. Prior work has largely focused on simple or global constraints—such as fixed boundaries or symmetries, that apply uniformly across the data distribution. Handling parameter-dependent constraints naively would require training over the joint distribution of solutions and parameters, which is often infeasible because parametric labels are missing, expensive to obtain, or high-dimensional. Addressing this limitation is critical for scientific discovery. Many inverse problems in the natural sciences and engineering require reasoning about unobserved parameters or exploring hypothetical scenarios inaccessible to direct experimentation. A generative model that can enforce parameter-dependent PDE constraints using only observational data would provide a powerful tool for data-efficient simulation, hypothesis testing, and the discovery of new physical phenomena, helping to bridge the gap between raw observations and mechanistic understanding.
@@
This work proposes a framework for fine-tuning flow-matching generative models to enforce parameter-dependent PDE constraints without requiring joint parameter–solution training data. This work aligns with a growing trend of simulation-augmented machine learning (Karniadakis et al., 2021), where generative models accelerate scientific discovery by efficiently exploring physically plausible solution spaces. Our approach reformulates fine-tuning as a stochastic optimal control problem via Adjoint Matching (Domingo-Enrich et al., 2025), guided by weak-form PDE residuals. By augmenting the model with a latent parameter evolution, we enable joint generation of physically consistent solution–parameter pairs, addressing ill-posed inverse problems. We evaluate our proposed fine-tuning framework on four representative PDE families spanning elliptic diffusion, elasticity, wave propagation, and incompressible flow and show an application to natural images. We demonstrate denoising and conditional generation capabilities, including robustness to noisy data and the ability to infer latent parameters from sparse observations. Visual and quantitative results, including strong reductions in residuals across tasks and robustness to model misspecification, highlight the flexibility of our method for integrating physical constraints into generative modelling.
@@ 2 Related Work
Integrating physical constraints—such as boundary conditions, symmetry invariances, and partial differential equation (PDE) constraints—into machine learning models improves both accuracy and out-of-distribution generalization. Classical approaches, such as Physics-Informed Neural Networks (PINNs, Raissi et al. (2019)), directly regress solutions that satisfy governing equations. While effective for forward or inverse problems, PINNs do not capture distributions over solutions, making them unsuitable for generative tasks that require sampling diverse plausible outcomes.
In the generative setting, the main challenge is ensuring that the physically constrained samples retain the variability of the underlying generative model, avoiding pathalogical issues such as mode collapse. Bastek et al. (2024) proposes a unified framework for introducing physical constraints into Denoising Diffusion Probabilistic Models (DDPMs, Ho et al. (2020)) at pre-training time, by adding a first-principles physics-residual loss to the diffusion training objective. This loss penalises violations of governing PDEs (e.g. fluid dynamics equations) so that generated samples inherently satisfy physical laws. The method was empirically shown to reduce residual errors for individual samples significantly, while simultaneously acting as a regulariser against overfitting, thereby improving generalisation. To evaluate the physics-residual loss, one needs to compute the expected PDE residual of the final denoised sample conditioned on the current noisy state in the DDPM process. Accurately estimating this expectation requires generating multiple reverse-diffusion trajectories from the same noisy sample, which makes pre-training significantly more expensive. A common alternative is to use Tweedie’s formula to approximate the conditional expectation in a single pass, but this shortcut introduces bias, particularly in the final denoising steps.
Zhang and Zou (2025) proposes enforcing constraints through a post-hoc distillation stage, where a deterministic student model is trained from a vanilla diffusion model to generate samples in one-step, regularised by a PDE residual loss. In Wang et al. (2025) the authors introduce PhyDA, diffusion-based data assimilation framework that ensures reconstructions obey PDE-based dynamics, specifically for atmospheric science. An autoencoder is used to encode sparse observations into a structured latent prior for the diffusion model, which is trained with an additional physical residual loss.
@@
Various works have proposed approaches to enforce PDE constraints at inference time, often in combination with observational constraints, drawing connections to conditional diffusion models (Dhariwal and Nichol, 2021; Ho and Salimans, 2021). Huang et al. (2024) introduce guidance terms within the denoising update of a score-based diffusion model to steer the denoising process towards solutions which are both consistent with data and underlying PDEs. A related approach was considered by Xu et al. (2025), further introducing an adaptive constraint to mitigate instabilities in early diffusion steps. In Christopher et al. (2024), the authors recast the inference-time sampling of a diffusion process as a constrained optimization problem, each diffusion step is projected to satisfy user-defined constraints or physical principles. This allows strict enforcement of hard constraints (including convex and non-convex constraints, as well as ODE-based physical laws) on the generated data. Lu and Xu (2024) consider the setting where the base diffusion model is trained on cheap, low-fidelity simulations, leveraging a similar approach to generate down-scaled samples via projection.
@@
Flow-matching (FM, Lipman et al. (2023)) has emerged as a flexible generative modelling paradigm for complex physical systems across science, including molecular systems (Hassan et al., 2024), weather (Price et al., 2023) and geology (Zhang et al., 2025) . In the context of physics-constrained generative models Utkarsh et al. (2025) introduces a zero-shot inference framework to enforce hard physical constraints in pretrained flow models, by repeatedly projecting the generative flow at sampling time. Similarly, Cheng et al. (2024) proposed the ECI algorithm, to adapt a pretrained flow-matching model so that it exactly satisfies constraints without using analytical gradients. In each iteration of flow sampling, ECI performs: an Extrapolation step (advancing along the learned flow), a Correction step (applying a constraint-enforcement operation), and an Interpolation step (adjusting back towards the model’s trajectory). While projection approaches are a compelling strategy for hard constraints, they can be challenging particularly for local constraints such as boundary conditions, as direct enforcement can introduce discontinuities. The above approach mitigates this by interleaving projections with flow steps, however this relies on the flow’s ability to rapidly correct such non-physical artifacts.
Baldan et al. (2025) propose Physics-Based Flow Matching (PBFM), which embeds constraints (PDE or symmetries) directly into the FM loss during training. The approach leverages temporal unrolling to refine noise-free final state predictions and jointly minimizes generative and physics-based losses without manual hyperparameter tuning of their tradeoff. To mitigate conflicts between physical constraints and the data loss, they employ the ConFIG (Liu et al., 2024), which combines the gradients of both losses in a way that ensures that gradient updates always minimise both losses simultaneously.
@@
Related to our approach are the works on generative models for Bayesian inverse problems (Stuart, 2010), where the goal is to infer distributions over latent PDE parameters given partial or noisy observations. Conditional diffusion and flow-matching models can be used to generate samples from conditional distributions and posterior distributions, supporting amortized inference and uncertainty quantification (Song et al., 2021; Utkarsh et al., 2025; Zhang et al., 2023). Conditioning is typically achieved either through explicit parameter inputs or guidance mechanisms during sampling, as in classifier-guided diffusion. While effective when large volumes of paired training data is available, these approaches are less relevant to observational settings where parameters are unobserved. In contrast, our approach connects the observed data to the latent parameters only during post-training, requiring substantially smaller volumes of data.
@@ 3 Method
FM models are trained to learn and sample from a given distribution of data $p_{\text{data}}$. They approximate this distribution by constructing a Markovian transformation from noise to data, such that the time marginals of this transformation match those of a reference flow $X_{t}=\beta_{t}X_{1}+\gamma_{t}X_{0}$. Specifically FM models learn a vector field $v_{t}(x)$ that transports noise to data, via the ODE $dX_{t}=v_{t}(X_{t})$. We can optionally inject a noise schedule $\sigma(t)$ along the trajectory to define an equivalent SDE that preserves the same time marginals (Maoutsa et al., 2020),
$dX_{t}=\left(v_{t}(X_{t})+\frac{\sigma(t)^{2}}{2\eta_{t}}\left(v_{t}(X_{t})-\frac{\dot{\beta}_{t}}{\beta_{t}}X_{t}\right)\right)+\sigma(t)\,dB_{t}\\ \ =:\ b_{t}(X_{t})+\sigma(t)\,dB_{t},$ (1)
where we combine coefficients $\beta_{t}$ and $\gamma_{t}$ into $\eta_{t}=\gamma_{t}\bigl(\tfrac{\dot{\beta}_{t}}{{\beta}}\gamma_{t}-\dot{\gamma}_{t}\bigr)$.
@@
To achieve this, we leverage the adjoint-matching framework of Domingo-Enrich et al. (2025). This work reformulates reward fine-tuning for flow-based generative models as a control problem in which the base generative process given by $v_{t}^{\text{base}}$ is steered toward high-reward samples via modifying the learned vector field, which we denote as $v_{t}^{\text{ft}}$ with corresponding drift term $b_{t}^{\text{ft}}$. Our approach is conceptually related to reward- or preference-based fine-tuning of generative models (Christiano et al., 2017; Sun et al., 2024), where a learned or computed reward steers generation toward desired properties. Here, the reward is defined via PDE residuals, encoding knowledge about underlying dynamics and physical constraints to the solutions space as deviations to differential operators or boundary conditions.
@@
3.1 Reward
A generative model can reproduce the visual characteristics of empirical data while ignoring the physics that governs it, thereby rendering the samples unusable for downstream scientific tasks. To bridge this gap we impose the known governing equations as soft constraints, expressed through differential operators $\mathcal{L}_{\alpha}x=0$ with parameters $\alpha$. Throughout, a generated sample $x$ is interpreted as the discretisation of a continuous field $x(\xi)$ on a domain $\Omega$. The strong PDE residual is defined as
$\mathcal{R}_{\text{strong}}(x,\alpha)=\bigl\|\mathcal{L}_{\alpha}x\bigr\|_{L^{2}(\Omega)}^{2}.$
In practice, strong residuals involve high-order derivatives that make the optimization landscape unstable. We therefore adopt weak-form residuals of the form $\langle\mathcal{L}_{\alpha}x,\psi\rangle_{L^{2}(\Omega)}$ for suitably chosen test functions $\psi\in\Psi$, which are numerically more stable under noisy or misspecified data. Repeated applications of integration-by-parts can transfer derivatives from $x$ to $\psi$. The set $\Psi$ is composed of compactly supported local polynomial kernels. For each evaluation we draw $N_{\text{test}}$ such functions; their centers and length‑scales are sampled at random. A mollifier enforces $\psi|_{\partial\Omega}=0$, justifying the integration by parts. The resulting residual is
@@ 3.1 Reward
These randomly sampled local test functions act as stochastic probes of PDE violations, providing a low-variance, data-efficient learning signal. A more detailed description of the test functions used can be found in Appendix D.3. Note that the residual might be augmented by adding soft constraints for boundary conditions.
3.2 Joint Evolution
@@
$\hat{x}_{1}\;=\;x_{t}+(1-t)\,v_{t}^{\text{base}}(x_{t}),\qquad\hat{\alpha}_{1}\;=\;\varphi(\hat{x}_{1}).$
The direction from the current state $\alpha_{t}$ to the predicted final parameter $\hat{\alpha}_{1}$ serves as a base vector field which we use to evolve alpha, i.e. $v_{t,\,\alpha}^{\text{base}}(\alpha_{t})=(\hat{\alpha}_{1}-\alpha_{t})/(1-t)$ inducing corresponding drift $b_{t,\alpha}^{\text{base}}$. This surrogate base flow, starting at a noise sample $\alpha_{0}^{\text{base}}\sim\mathcal{N}(0,I)$, emulates a denoising process of the recovered parameter. We denote by $\alpha^{\text{base}}$ the parameter aligned with the base trajectory $x^{\text{base}}$. While the evolution of $\alpha^{\text{base}}$ does not influence the trajectory of $x^{\text{base}}$, the inferred vector field can be used to effectively regularize the generation of the fine-tuned model. Similarly, to regularise towards the parameter recovered under the base model, we introduce an additional field $v_{t,\,\alpha}^{\text{reg}}(\alpha_{t}^{\text{ft}})=(\hat{\alpha}_{1}^{\text{base}}-\alpha_{t}^{\text{ft}})/(1-t)$. This vector field points from the current parameter estimate of the fine-tuned trajectory $\alpha_{t}^{\text{ft}}$ to the recovered parameter under the base model $\hat{\alpha}_{1}^{\text{base}}$. The field is used to pull the fine‑tuned dynamics towards final samples associated with parameters similar to those of the base trajectory. The introduced vector fields are visualized in Fig. 1.
3.3 Adjoint Matching